5. An unfingered guitar string is 0.70 m long and is tuned to play E (330 Hz). The first harmonic.
a. What is the speed of the wave on the string?
b. How far from the end of this string must the ginger be placed to play an A (440 Hz)?
1
Expert's answer
2020-04-23T12:58:40-0400

a.

"v=2lf=2(0.7)(330)=462\frac{m}{s}"

b.

"2(0.7-x)(440)=462"

"x=0.175\ m"